1/A batch contains 10 transistors of which three are defective. If three are chosen, in how many ways can one get two defective?
2/ Three marbles are chosen from an urn that contains 5 reds, 4 white, and 3 blue marbles. how many samples if we choose all three of the same color?
I very appreciate your help !!!

Answer by jim_thompson5910(35256)

I'll do the first one to get you started.

# 1

10 transistors
3 defective
7 nondefective


Number of ways to get exactly 2 defective

(3 C 2)*(7 C 1) = (3)*(7) = 21

There are 21 ways to pick exactly 2 defective transistors (out of 3 selected transistors).

Side note: I'm using the combination formula . Many calculators such as the TI84 have this formula built in (let me know if you need steps on how to use it).

Another side note: 3 C 2 represents picking exactly 2 bad transistors out of 3. 7 C 1 represents picking exactly 1 good transistor out of 7.